# Logical Consistency of Large Language Models in Fact-Checking

### v1

@@ Appendix A Limitations and Broader Impacts
LLM’s general reasoning ability has been under much scrutiny in recent years, and experiments suggest that there is still a gap between LLMs and human-like logical reasoning ability. Given such a state of affairs, one limitation of our current study is that we do not evaluate how much logical consistency improves the general reasoning ability of LLMs. Another limitation is that, our paper mainly focuses on binary responses, and do not consider extensions to multi-class logic and non-binary answer consistency. We will consider such extension in the future.
Appendix B Related Work
@@ B.1 Large Language Models with KG Contexts
Knowledge graphs (KGs) can enhance pre-training (Yu et al., 2022), fine-tuning (Wang et al., 2024b), inference (Wei et al., 2023), prompting (Baek et al., 2023), retrieval/ knowledge augmented generation (Xu et al., 2024), knowledge editing (Zheng et al., 2023), and knowledge validation (Hu et al., 2024a) of LLMs. In particular, KGs can assist in the inference phase of LLMs through retrieval-augmented generation (RAG), e.g., to provide KG context within prompts for reducing hallucinations, thus improving accuracy and consistency. For instance, Liu et al. (2020) inject KG triplets into LLMs to enhance the domain knowledge, meanwhile Li et al. (2023) use external data sources to improve the factual correctness of large language models. GraphRAG is the latest technology that models an external knowledge base as a knowledge graph to improve retrieval methods with a more comprehensive contextual understanding and thereby assists in obtaining more precise search results via LLMs (Xu et al., 2024; Edge et al., 2024; Sanmartin, 2024; Liu et al., 2024). Since we consider a RAG paradigm for LLM’s consistency assessment, a good retrieval technique (e.g., future developments on graphRAG) can be leveraged in our framework to further improve the LLM’s accuracy and consistency in fact-checking.
@@ B.2 Measuring and Improving the Consistency of LLMs
With the prevalence of large language models, measuring their consistency has become critical. LLMs can generate logically contradicting outputs, e.g., different predictions for meaning-preserving text alternations (Ravichander et al., 2020; Elazar et al., 2021). They may violate important relational properties such as negation (Kassner & Schütze, 2020; Ettinger, 2020; Asai & Hajishirzi, 2020; Jang et al., 2022), symmetry (Wang et al., 2019; Li et al., 2019; Kumar & Joshi, 2022), and transitivity (Asai & Hajishirzi, 2020; Lin & Ng, 2022). Raj et al. (2023) measure the semantic similarity of LLM outputs in response to paraphrased versions of a question.
@@
First, unlike bulk of the literature that measure an LLM’s consistency on natural language processing tasks, we develop systematic consistency measurement techniques of retrieval-augmented LLMs for propositional logic facts and logic rules, with knowledge graph contexts. Second, existing approaches generally update all model parameters and are inefficient. We avoid full fine-tuning by using parameter-efficient supervised fine-tuning through low rank adaption. We have experimentally demonstrated that our approach is both effective, efficient, and generalizes to complex facts and logic rules, large-scale KGs, and LLMs having billions of parameters.
@@ B.3 Large Language Models Hallucinations
LLMs may not explicitly store a consistent representation of knowledge. Hence, they can hallucinate by generating unreliable and incoherent responses, and often plausible yet factually incorrect statements. There are several reasons and categories of LLM hallucinations. In this work, our focus is faithfulness hallucination (Huang et al., 2024a), resulting in divergence of LLM generated contents from user inputs (e.g., KG contexts, instructions, examples, etc.), as well as the lack of consistency within the generated contents when the input queries are logically perturbed. Subsequently, various strategies have been developed to detect and reduce LLMs’ hallucinations. For instance, retrieval-augmented generation (RAG) enhances models with up-to-date knowledge from external sources. More advanced approaches (Wang et al., 2024a) retrieve high-quality candidates to avoid the context poisoning issue, e.g., LLM-Embedder (Lv et al., 2024) is a general-purpose retrieval technique that learns vector embedding by rewarding high-quality retrieval candidates, contrastive learning, and knowledge distillation. In fact, we have already shown that vector embedding methods improve logical consistency over traversal-based context retrieval (Appendix E, Table 11). In future, it would be interesting to replace the Sentence Transformer in our vector embedding method via LLM-Embedder to construct more relevant contexts. Another approach to reduce LLMs’ hallucinations is the chain-of-thought (CoT) prompting (Wei et al., 2022), which encourages LLMs to generate reasoning steps before responding. We have also conducted experiments with CoT (Appendix E, Table 12), but it does not demonstrate any improvement over fine-tuning. As such, we hypothesise that achieving logical consistency requires an update in the internal weights of the LLMs, e.g., via supervised fine-tuning. Along the same direction, self-consistency (Wang et al., 2023a) employs a post-processing approach that does not improve the internal weights of the model and instead addresses the decoding module of the model – it samples multiple responses of the LLM and chooses a response with the majority vote that is denoted as consistent. In contrast, our fine-tuning procedure improves the internal weights of the LLM to be a consistent responder under logical perturbation of input queries, thereby reducing faithfulness hallucination.
@@ B.4 Fact-checking with KGs and LLMs
Knowledge graphs, reliably storing facts data, are valuable sources in fact-checking. We refer to Luo & Long (2021); Huynh & Papotti (2018) for surveys and experimental benchmarks. In particular, KG approaches for fact-checking can be broadly classified into paths (Shi & Weninger, 2016), rules and patterns (Lin et al., 2019; Galárraga et al., 2015), and embedding-based (Ammar & Celebi, 2019; Borrego et al., 2021). Recent embedding methods determine whether a given fact, currently missing in an incomplete knowledge graph, should appear in it. In contrast, our focus is to validate logic facts from the input KG. LLMs like GPT-3.5 and GPT-4 have also been exploited for automated fact-checking with and without access to external contexts, e.g., Google search results (Quelle & Bovet, 2024). Hamed et al. (2023) studied fact-checking in biological graphs constructed from ChatGPT contents. Recently, MiniCheck (Tang et al., 2024) and RefChecker (Hu et al., 2024b) perform fact-checking of LLM outputs, primarily simple facts or claim-triplets, on grounding documents. To the best of our knowledge, ours is the first work on measuring and improving the consistency of LLMs for propositional logic-based fact-checking with KG contexts.
Appendix C Proofs of Propositions 1 and 2
@@ Appendix D Additional Discussions on Methodology
Supervised Fine-tuning. We adopted a parameter-efficient fine-tuning (PEFT) method based on QLoRA (Quantized Low-Rank Adaptation) (Dettmers et al., 2024), which is a more memory-efficient adaptation of LoRA (Hu et al., 2021a). The main idea behind LoRA is to freeze the pre-trained model weight-matrix $W_{0}$ of size $d\times d$ and represent the trainable update-matrix ($\Delta W$) as two low-rank matrices $A$ and $B$ of dimensions $d\times r$ and $r\times d$, respectively. Here, $r<<d$, e.g., $r=64$ and $d\approx 16$K for Llama2222-$7$B and $d\approx 8$K for Gemma-$2$B in our experiments. Given an input vector $x$, during forward-pass, instead of computing $(W_{0}+\Delta W)x$, LoRA computes $(W_{0}x+BAx)$. The benefit of storing two low-rank matrices instead of $\Delta W$ is that the model needs to update only $2rd$ parameters, which is more efficient than updating $d^{2}$ parameters. QLoRA quantizes the precision of the weights to 4 bits. Due to quantization, QLoRA is more memory efficient; for instance, it has been reported that one can fine-tune a 65B parameter Llama model on a single 48GB GPU, while preserving full 16-bit finetuning task performance (Dettmers et al., 2024).
Appendix E Additional Experiments
Datasets.
@@ Appendix E Additional Experiments
More complex and higher-order logic datasets. We have curated two more datasets from Freebase containing more complex logical expressions, especially the law of Syllogism (LoS) and the first-order logic (FoL). In the LoS dataset, there are 100 queries of the form $r_{1}(p,q)\land r_{2}(q,s)$ encapsulating the logic $(p\Rightarrow q)\land(q\Rightarrow s)$, which, according to the law of Syllogism should be equivalent to $p\Rightarrow s$. An example: $Contains(Oceania,NewZealand)\land Country(NewZealand,Auckland)$ is logically equivalent to $Contains(Oceania,Auckland)$.
In the FoL dataset, there are 500 queries of the form $\exists_{x}r_{1}(p,x)\land r_{2}(x,q)$ and their negations in the form of $\forall_{x}\neg r_{1}(p,x)\lor\neg r_{2}(x,q)$. The negation of the former should be logically equivalent to the latter.
Parameters. For an efficient fine-tuning, we adopted QLoRA implementation from Huggingface (Dettmers et al., 2024). In Llama2222-$7$B and $13$B, the hyperparameter choices of QLoRA are the following: learning rate $2\times 10^{-5}$, weight decay $0.001$, warm-up ratio $0.03$, batch size $8$, $r=16$, $\alpha=16$, and dropout $0.1$. In Gemma-$2$B, we consider learning rate $2\times 10^{-6}$ and keep other hyper-parameters similar to Llama. For high-throughput and memory-efficient inference, we adopt the vLLM library (Kwon et al., 2023), and set temperature to 00 for a deterministic output. We limit the context length in the LLMQuery by selecting relevant triplets of around $1000$ tokens.
System Setup. All experiments are conducted in Python version $3.8.0$. Fine-tuning is conducted on a cluster with two NVIDIA A40404040 $45$ GB GPUs having Intel(R) Xeon(R) Gold $5317$ CPU @ $3.00$ GHz, $48$ core and $1007$ GB RAM. Inference is conducted on a cluster with two Tesla V100100100100-PCIE $32$ GB GPUs having Intel(R) Xeon(R) Gold $6134$M CPU @ $3.20$ GHz, $32$ core, and $755$ GB RAM.
@@
Efficiency. We compare the per-epoch running time of PEFT vs. full fine-tuning of Llama2222-$7$B (Table 6). PEFT requires on average $4.43$ hours to complete an epoch including the computation of evaluation loss, whereas full fine-tuning requires $12.52$ hours. Therefore, PEFT is more efficient than full fine-tuning in improving the efficiency of supervised fine-tuning for logical consistency.
@@
Dataset Fact BFS Context Create LLM Inference Total
@@
Llama2222 $7$B FreebaseLFC $0.51$ $\mathbf{0.88}$ $0.03$ $\mathbf{0.77}$
Impact of KG Contexts in LLMQuery. Table 8 shows the effectiveness of using a context on the accuracy and logical consistency of Llama2222-$7$B for simple facts. Adding a context provides the LLM a premise to answer the subsequent fact. As such, Llama2222-$7$B base model demonstrates a higher accuracy and logical consistency in the presence of a context, and hence, in all our experiments we have provided a KG context to the LLMQuery.
Table 9: Few-shot prompting results on Llama2222-$7$B model.
Accuracy Logical Consistency
Dataset Before FT 0-shot Few-shot After FT Before FT 0-shot Few-shot After FT
FreebaseLFC $0.90$ $0.89$ 0.69 $\mathbf{0.93}$ $0.81$ 0.79 0.40 $\mathbf{0.86}$
Impact of Context Length.
@@
Impact of Different Context Retrieval Methods.
@@
Dynamically Varying Contexts. Real-world applications may need context retrieval to quickly adapt to new information, e.g., the knowledge graph evolving with time. In the case of evolving KGs, the “BFS + Relation hierarchy” retrieval is a computationally better alternative than embedding-based retrieval methods such as “BFS + Vector Embedding” or “Vector Embedding (without BFS)” method. Although the embedding-based methods slightly improve the accuracy and consistency (as shown in Table 11), due to their higher pre-processing overhead, they must be conducted offline repeatedly as the KG changes to obtain better entity embeddings.
Table 12: Few-shot and Chain-of-Thought (CoT) prompting results on Llama2222-$7$B model.

### v2

@@ Appendix A Limitations and Broader Impacts
LLM’s general reasoning ability has been under much scrutiny in recent years, and experiments suggest that there is still a gap between LLMs and human-like logical reasoning ability. Given such a state of affairs, one limitation of our current study is that we do not evaluate how much logical consistency improves the general reasoning ability of LLMs. Another limitation is that our paper mainly focuses on binary responses, and does not consider extensions to multi-class logic and non-binary answer consistency. We will consider such extensions in the future.
Appendix B Related Works
@@ B.1 Large Language Models with KG Contexts
Knowledge graphs (KGs) can enhance pre-training (Yu et al., 2022), fine-tuning (Wang et al., 2024b), inference (Wei et al., 2023), prompting (Baek et al., 2023), retrieval/ knowledge augmented generation (Xu et al., 2024), knowledge editing (Zheng et al., 2023), and knowledge validation (Hu et al., 2024a) of LLMs. In particular, KGs can assist in the inference phase of LLMs through retrieval-augmented generation (RAG), e.g., to provide KG context within prompts for reducing hallucinations, thus improving accuracy and consistency. For instance, Liu et al. (2020) inject KG triplets into LLMs to enhance the domain knowledge, meanwhile Li et al. (2023) uses external data sources to improve the factual correctness of large language models. GraphRAG is the latest technology that models an external knowledge base as a knowledge graph to improve retrieval methods with a more comprehensive contextual understanding and thereby assists in obtaining more precise search results via LLMs (Xu et al., 2024; Edge et al., 2024; Sanmartin, 2024; Liu et al., 2024). Since we consider a RAG paradigm for LLM’s consistency assessment, a good retrieval technique (e.g., future developments on graphRAG) can be leveraged in our framework to further improve the LLM’s accuracy and consistency in fact-checking.
@@ B.2 Measuring and Improving the Consistency of LLMs
With the prevalence of large language models, measuring their consistency has become critical. LLMs can generate logically contradicting outputs, e.g., different predictions for meaning-preserving text alternations (Ravichander et al., 2020; Elazar et al., 2021). They may violate important relational properties such as negation (Kassner & Schütze, 2020; Ettinger, 2020; Asai & Hajishirzi, 2020; Jang et al., 2022), symmetry (Wang et al., 2019; Li et al., 2019; Kumar & Joshi, 2022), and transitivity (Asai & Hajishirzi, 2020; Lin & Ng, 2022). Raj et al. (2023) measures the semantic similarity of LLM outputs in response to paraphrased versions of a question.
@@
First, unlike the bulk of the literature that measures an LLM’s consistency on natural language processing tasks, we develop systematic consistency measurement techniques of retrieval-augmented LLMs for propositional logic facts and logic rules, with knowledge graph contexts. Second, existing approaches generally update all model parameters and are inefficient. We avoid full fine-tuning by using parameter-efficient supervised fine-tuning through low-rank adaption. We have experimentally demonstrated that our approach is effective, efficient, and generalizes to complex facts and logic rules, can handle large-scale KGs, and LLMs having billions of parameters.
@@ B.3 Large Language Models Hallucinations
LLMs may not explicitly store a consistent representation of knowledge. Hence, they can hallucinate by generating unreliable and incoherent responses, and often plausible yet factually incorrect statements. There are several types of LLM hallucinations. In this work, our focus is faithfulness hallucination (Huang et al., 2024a), resulting in divergence of LLM-generated contents from user inputs (e.g., KG contexts, instructions, examples, etc.), as well as the lack of consistency within the generated contents when the input queries are logically perturbed. Subsequently, various strategies have been developed to detect and reduce LLMs’ hallucinations. For instance, retrieval-augmented generation (RAG) enhances models with up-to-date knowledge from external sources. More advanced approaches (Wang et al., 2024a) retrieve high-quality candidates to avoid the context poisoning issue, e.g., LLM-Embedder (Lv et al., 2024) is a general-purpose retrieval technique that learns vector embedding by rewarding high-quality retrieval candidates, contrastive learning, and knowledge distillation. In fact, we have already shown that vector embedding methods improve logical consistency over traversal-based context retrieval (Appendix E, Table 11). In the future, it would be interesting to replace the Sentence Transformer in our vector embedding method via LLM-Embedder to construct more relevant contexts. Another approach to reduce LLMs’ hallucinations is the chain-of-thought (CoT) prompting (Wei et al., 2022), which encourages LLMs to generate reasoning steps before responding. We have also conducted experiments with CoT (Appendix § G.5.3), but it does not demonstrate any improvement over fine-tuning. As such, we hypothesize that achieving logical consistency requires an update in the internal weights of the LLMs, e.g., via supervised fine-tuning. Moreover, self-consistency (Wang et al., 2023a) employs a post-processing approach that does not improve the internal weights of the model and instead addresses the decoding module of the model – it samples multiple responses of the LLM and chooses a response with the majority vote that is denoted as consistent. In contrast, our fine-tuning procedure improves the internal weights of the LLM to be a consistent responder under logical perturbation of input queries, thereby reducing faithfulness hallucination.
@@ B.4 Fact-checking with KGs and LLMs
Knowledge graphs that store high-quality facts are reliable and valuable sources for fact-checking. We refer to Luo & Long (2021); Huynh & Papotti (2018) for surveys and experimental benchmarks. In particular, KG approaches for fact-checking can be broadly classified into paths (Shi & Weninger, 2016), rules and patterns (Lin et al., 2019; Galárraga et al., 2015), and embedding-based (Ammar & Celebi, 2019; Borrego et al., 2021). Recent embedding methods determine whether a given fact, currently missing in an incomplete knowledge graph, should appear in it. In contrast, our focus is to validate logic facts from the input KG.
LLMs like GPT-3.5 and GPT-4 have also been exploited for automated fact-checking with and without access to external contexts, e.g., Google search results (Quelle & Bovet, 2024)Hamed et al. (2023) studied fact-checking in biological graphs constructed from ChatGPT contents. Recently, MiniCheck (Tang et al., 2024) and RefChecker (Hu et al., 2024b) perform fact-checking of LLM outputs, primarily simple facts or claim-triplets, on grounding documents. To the best of our knowledge, ours is the first work on measuring and improving the consistency of LLMs for propositional logic-based fact-checking with KG contexts.
Appendix C Formal Proofs
@@
Proposition 3.
An LLM is consistent with respect to a CNF fact $q=\wedge_{i=1}^{n}c_{i}$, where $c_{i}=\vee_{j=1}^{i_{m}}e_{ij}$, if
$\mathtt{LLM}(q)=\bigwedge_{i=1}^{n}\left(\bigvee_{j=1}^{i_{m}}\mathtt{LLM}(e_{ij})\right).$
Proof.
Applying the definition of consistency to the LLM response to $q$, we obtain
$\mathtt{LLM}(q)$ $=\mathtt{LLM}(c_{1})\wedge\mathtt{LLM}(c_{2})\wedge\ldots\wedge\mathtt{LLM}(c_{n})$
$=\mathtt{LLM}(\bigvee_{j=1}^{1_{m}}e_{1j})\wedge\mathtt{LLM}(\bigvee_{j=1}^{2_{m}}e_{2j})\wedge\ldots\wedge\mathtt{LLM}(\bigvee_{j=1}^{n_{m}}e_{nj})$
$=\left(\bigvee_{j=1}^{1_{m}}\mathtt{LLM}(e_{1j})\right)\wedge\left(\bigvee_{j=1}^{2_{m}}\mathtt{LLM}(e_{2j})\right)\wedge\ldots\wedge\left(\bigvee_{j=1}^{n_{m}}\mathtt{LLM}(e_{nj})\right)$
$=\bigwedge_{i=1}^{n}\left(\bigvee_{j=1}^{i_{m}}\mathtt{LLM}(e_{ij})\right)$
@@ Appendix D Additional Discussions on Methodology
D.1 Supervised Fine-tuning
We adopted a parameter-efficient fine-tuning (PEFT) method based on QLoRA (Quantized Low-Rank Adaptation) (Dettmers et al., 2024), which is a more memory-efficient adaptation of LoRA (Hu et al., 2021a). The main idea behind LoRA is to freeze the pre-trained model weight-matrix $W_{0}$ of size $d\times d$ and represent the trainable update-matrix ($\Delta W$) as two low-rank matrices $A$ and $B$ of dimensions $d\times r$ and $r\times d$, respectively. Here, $r<<d$, e.g., $r=64$ and $d\approx 16$K for Llama2222-$7$B and $d\approx 8$K for Gemma-$2$B in our experiments. Given an input vector $x$, during forward-pass, instead of computing $(W_{0}+\Delta W)x$, LoRA computes $(W_{0}x+BAx)$. The benefit of storing two low-rank matrices instead of $\Delta W$ is that the model needs to update only $2rd$ parameters, which is more efficient than updating $d^{2}$ parameters. QLoRA quantizes the precision of the weights to 4 bits. Due to quantization, QLoRA is more memory efficient; for instance, it has been reported that one can fine-tune a 65B parameter Llama model on a single 48GB GPU while preserving full 16-bit finetuning task performance (Dettmers et al., 2024).
Appendix E Datasets
E.1 Data Sources
We consider existing KGs, such as Freebase, NELL, and WikiKG90Mv2, and convert to logical fact-checking datasets. Freebase contains real-world entities such as people, places, institutions, and their relations from the Freebase KG. NELL is a dataset built from the Web via an intelligent agent called Never-Ending Language Learner. This agent attempts to learn over time to read the web. NELL has accumulated over 100 million candidate triplets by reading the web (April 2018) with different levels of confidence. WikiKG90Mv2 is a KG extracted from the entire Wikidata knowledge base.
@@ E.1 Data Sources
E.2 Our curated datasets
[…]
Figure 3: (Continuing Figure 1) A representative $\mathtt{LLMQuery}$ on a simple fact used in our experiments. The expected LLM response is in blue color.
Consider the context: \$Context.Is the fact correct given the context? \$Fact-1 and \$Fact-2.$\mathtt{LLMQuery}$ on (\$Fact-1$\wedge$\$Fact-2, \$Context)Consider the context: \$Context.Is the fact correct given the context? \$Fact-1 or \$Fact-2.$\mathtt{LLMQuery}$ on (\$Fact-1$\vee$\$Fact-2, \$Context)
Figure 4: (Continuing Figure 1) $\mathtt{LLMQuery}$ when considering conjunctive and disjunctive facts. We replace logical operators with their natural language description, such as $\wedge$ with ‘and’ and $\vee$ with ‘or’, when constructing $\mathtt{LLMQuery}$.
@@ E.2 Our curated datasets
E.3 More complex and higher-order logic datasets
We have curated two more datasets from Freebase containing more complex logical expressions, especially the law of Syllogism (LoS) and the first-order logic (FoL). In the LoS dataset, there are 100 queries of the form $r_{1}(p,q)\land r_{2}(q,s)$ encapsulating the logic $(p\Rightarrow q)\land(q\Rightarrow s)$, which, according to the law of Syllogism should be equivalent to $p\Rightarrow s$.
[…]
In the FoL dataset, there are 500 queries of the form $\exists_{x}r_{1}(p,x)\land r_{2}(x,q)$ and their negations in the form of $\forall_{x}\neg r_{1}(p,x)\lor\neg r_{2}(x,q)$. The negation of the former should be logically equivalent to the latter. In $\mathtt{LLMQuery}$, we consider the natural language description ‘There exists an entity $x$ such that’ and ‘For all entity $x$ such that’ to express existential $\exists$ and for-all $\forall$ quantifiers, respectively.
Appendix F Parameter settings and System configurations
F.1 Parameter settings
For an efficient fine-tuning, we adopted QLoRA implementation from Huggingface (Dettmers et al., 2024). In Llama2222-$7$B and $13$B, the hyperparameter choices of QLoRA are the following: learning rate $2\times 10^{-5}$, weight decay $0.001$, warm-up ratio $0.03$, batch size $8$, $r=16$, $\alpha=16$, and dropout $0.1$. In Gemma-$2$B, we consider learning rate $2\times 10^{-6}$ and keep other hyper-parameters similar to Llama. For high-throughput and memory-efficient inference, we adopt the vLLM library (Kwon et al., 2023), and set temperature to 00 for a deterministic output. We limit the context length in the LLMQuery by selecting relevant triplets of around $1000$ tokens.
F.2 System Setup
All experiments are conducted in Python version $3.8.0$. Fine-tuning is conducted on a cluster with two NVIDIA A40404040 $45$ GB GPUs having Intel(R) Xeon(R) Gold $5317$ CPU @ $3.00$ GHz, $48$ core and $1007$ GB RAM. Inference is conducted on a cluster with two Tesla V100100100100-PCIE $32$ GB GPUs having Intel(R) Xeon(R) Gold $6134$M CPU @ $3.20$ GHz, $32$ core, and $755$ GB RAM.
Appendix G Additional Experiments
G.1 Efficiency
We compare the per-epoch running time of PEFT vs. full fine-tuning of Llama2222-$7$B (Table 6). PEFT requires on average $4.43$ hours to complete an epoch including the computation of evaluation loss, whereas full fine-tuning requires $12.52$ hours. Therefore, PEFT is more efficient than full fine-tuning in improving the efficiency of supervised fine-tuning for logical consistency.
@@ G.1 Efficiency
Dataset Fact BFS Context Creation LLM Inference Total
@@ G.2 Impact of KG Contexts
Llama2222-$7$B FreebaseLFC $0.51$ $\mathbf{0.88}$ $0.03$ $\mathbf{0.77}$
Table 8 shows the effectiveness of using a context on the accuracy and logical consistency of Llama2222-$7$B for simple facts. Adding a context provides the LLM a premise to answer the subsequent fact. As such, Llama2222-$7$B base model demonstrates a higher accuracy and logical consistency in the presence of a context, and hence, in all our experiments we have provided a KG context to the LLMQuery.
G.3 Impact of Context Length
@@
G.4 Impact of Different Context Retrieval Methods
@@
Handling Dynamically Varying Contexts.
Real-world applications may need context retrieval to quickly adapt to new information, e.g., the knowledge graph evolving with time. In the case of evolving KGs, the “BFS + Relation hierarchy” retrieval is a computationally better alternative than embedding-based retrieval methods such as “BFS + Vector Embedding” or “Vector Embedding (without BFS)” method. Although the embedding-based methods slightly improve the accuracy and consistency (as shown in Table 10), due to their higher pre-processing overhead, they must be conducted offline repeatedly as the KG changes to obtain better entity embeddings.
G.5 Results with Zero-Shot, 2-shot and Chain-of-thought Prompting
G.5.1 Supervised Fine-tuning vs Zero-shot Prompting
We demonstrate the extended results on comparing instruction prompting with fine-tuning in Table 11. We find that fine-tuning is superior to instruction prompting on simple facts, across datasets and models.
Table 11: Impact of instruction prompting (Prompt) vs. fine-tuning (FT) on simple facts. Fine-tuning results in higher accuracy and consistency than instruction prompting.